# Robust preparation of ground state phases under noisy imaginary time evolution

### v1

@@ I Introduction
Another conceptual approach to ground state preparation, which can potentially be more suitable for near-term implementation, is based on imaginary time evolution (ITE) [3, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Being a non-unitary operation, the ITE cannot be implemented using only unitary gates; nevertheless, various algorithms have been designed that effectively realize an approximate ITE process on a quantum device. These algorithms either (1) perform local tomography to determine a unitary such that the unitarily-evolved state approximates the ITE-evolved state [11, 12, 13, 14, 15, 16] (this is known as quantum imaginary time evolution, or QITE), (2) classically construct an adiabatic protocol that emulates the ITE [17] (this is known as adiabatic quantum imaginary time evolution, or A-QITE), (3) take advantage of the variational principle [18, 19, 20, 21, 22, 23, 24] (this is known as variational quantum imaginary time evolution, or VQITE), (4) utilize ancilla(e) and measurements [25, 26, 27] (this is known as probabilistic imaginary time evolution, or PITE), or (5) ancilla(e), measurements, and feedback [28]. These and other sophisticated protocols become increasingly realizable on present-day and near-term devices due to rapid hardware advances, such as the ever more efficient implementation of mid-circuit measurements and the adoption of dynamical circuits [29, 30, 31].
@@
In this work we ask whether ITE-based ground state preparation is robust to various types of noise commonly present in experimental setups. To keep our considerations as general and as simple as possible, we refrain from directly modeling a particular algorithm implementing the ITE, and instead consider a black box realizing a Trotterized version of ITE  [^1] This Trotterized approach to ITE is also implemented directly in a variety of classical algorithms for quantum many-body physics, including time-evolving block decimation [104, 105] and various quantum Monte Carlo approaches [106, 107, 108, 109]. A “circuit” constructed in this way can be thought of as a toy model of an ITE protocol, which we further subject to noise (see Fig. 1). Within this model the ITE effectively realizes an approximate (due to the Trotter error) weak projection onto a target ground state, which may belong to a nontrivial phase, whereas noise tends to trivialize the state it acts upon. Here we analyze this competition between ITE and noise by investigating whether the mixed state resulting from the noisy ITE process belongs to the same nontrivial phase as the targeted pure ground state. Note that this is different from the conventional task of the quantum state preparation, which amounts to preparing a state that maximizes the overlap with a target state.
@@
As a concrete example, we focus on the one-dimensional transverse-field Ising model whose ground-state quantum phase transition is a prototypical example of long-range order and spontaneous symmetry breaking (SSB). We investigate whether this order, and the quantum critical point from which it emerges, persists in the mixed steady state of the ITE at finite noise strength. To this end, in the limit of infinitesimal imaginary Trotter step we take advantage of the Choi-Jamiołkowski isomorphism [34, 35] and map the noisy ITE of interest to a noiseless ITE under an effective Hamiltonian in a doubled Hilbert space. This allows us to extract the steady state properties of the noisy ITE from the ground state of the effective Hamiltonian, which can be efficiently calculated using the density matrix renormalization group (DMRG) [36, 37, 38]. To further corroborate our findings and investigate the effect of imaginary-time Trotter error, we also perform direct simulations of our noisy ITE “circuit” for comparatively small system sizes. We find that the ordered phase is stable to sufficiently weak noise as long as the noise does not explicitly break the weak version [39, 40] of the symmetry that is spontaneously broken at the ground-state quantum critical point. These findings are in contrast to the naive expectation that noise should act similarly to a finite temperature, and thus should destroy any order and criticality in one dimension.
Our study connects to investigations of phases and phase transitions in open quantum systems [41, 42, 43, 44]. A growing body of research has focused on defining, characterizing, and designing mixed-state versions of symmetry protected topological (SPT) phases, quantum critical states, and topological order [45, 40, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79]. Many of these works specifically describe the fate of these nontrivial phases and states under noisy channels (see, for example, Refs. [51, 52, 50]) and employ the doubled Hilbert space formalism in their analyses of mixed states [52, 53, 51]. The present work differs in that the ITE tends to counteract the trivializing effect of the noise—heuristically, if the ITE superoperator is gapped, the added noise acts as a perturbation that can only destroy an ordered phase by closing the gap. This brings additional connections to passive error correction [80, 81, 59] and adaptive quantum circuits [82, 83, 84, 85].
@@ III.1.1 Single-qubit bit-flip noise
where in the second equality we have used the fact that $e^{\mu\sum_{i}\sigma^{x}_{i}\otimes\sigma^{x}_{i}}=\prod_{i}e^{\mu\sigma^{x}_{i}\otimes\sigma^{x}_{i}}=\prod_{i}[\cosh\mu+\sinh\mu\,\sigma^{x}_{i}\otimes\sigma^{x}_{i}]$ and have introduced a parameter $\mu=\tanh^{-1}[p/(1-p)]$. This parameter is real and positive for $20<p<1/$ 2 and in the limit of weak noise, $p\ll 1$, becomes $\mu\sim p+O(p^{2})$. Given Eqs. (7) and (12), in the doubled space the steady state of the noisy ITE can be written, up to an overall normalization prefactor, as
@@
To showcase the above statements, we numerically investigate [89] the symmetry breaking order-disorder transition in the steady state of the noisy ITE by computing the ground state of the spin ladder Hamiltonian (16)-(17) using the DMRG [36, 37, 38] and then calculating the expectation values of observables in the steady state via Eq. (9). To identify the location of the transition as precisely as possible, we calculate the Binder cumulant [90, 91],
@@
Figure 3(a) depicts the Binder cumulant evaluated for the DMRG-calculated ground state of the spin ladder Hamiltonian (16)–(17) as a function of the transverse field $g$ for various values of the ladder length $L$ and open boundary conditions. The Ising couplings along the legs of the ladder are taken to be ferromagnetic, while the strength of the noise-induced couplings along the rungs is set to $\lambda=0.1$. For this value of $\lambda$ the critical point, identified through the scale invariance of the Binder cumulant, is located at $g_{c}\approx 0.94$, which is slightly lower than the noiseless value of $g_{c}=1$. The ordered phase is thus still present in the noisy model, although its extent in parameter space is smaller when compared to the noiseless case. The curves representing different system sizes collapse when the Binder cumulant is plotted as a function of $(g-g_{c})L$, which is illustrated by the inset in Fig. 3(a). The finite size scaling thus yields the correlation length critical exponent of $\nu=1$, which, as expected, belongs to the 2D Ising universality class.
@@
We compute the two ground states, one with $\lambda=0.1$ and the other with $\lambda=0$, and the overlap between them using DMRG, and plot the resulting fidelity as a function of the transverse field in Fig. 4. For $g\ll 1$ ($g\gg 1$), both the noisy and the noiseless steady state are deep in their respective ordered (disordered) phases, and the fidelity between them is close to one. On the other hand, for values of $g$ in between the noiseless and the noisy critical points (marked by dotted lines in the figure), the noisy steady state is disordered, while the noiseless one is ordered, and the fidelity dips drastically toward zero. The dip is stronger for larger system sizes, with the fidelity reaching zero for $8002L=$ 800. We note that the fidelity also decreases with system size even deep in the ordered/disordered phase, albeit the decrease is much smaller there than in the region between the critical points.
@@ III.2 Amplitude damping noise channel
Figure 7: Binder cumulant plotted as a function of the transverse field and calculated using the DMRG for (a) the ferromagnetic (FM, $J=+1$) and (b) the antiferromagnetic (AFM, $J=-1$) Ising spin ladder model (16) when the interleg coupling is induced by the amplitude daming (AD) noise given by Eq. (32) with $\lambda_{z}=\lambda_{+}=0.4$. The curves are plotted for various ladder lengths $L$ in order to identify the absence (in the FM case) or the presence (in the AFM case) of the scale invariant critical point $g_{c}$; the total number of sites in the DMRG calcualtion is equal to $2L$. The inset in (b) depicts the Binder cumulant as a fucntion of $(g-g_{c})L$ and demonstrates the collapse of the curves corresponding to different system sizes. Open boundary conditions are assumed throughout the calculations in both (a) and (b).
@@
If the Ising ladder has ferromagnetic couplings along its legs, the Hamiltonian (32), in addition to breaking the strong $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry, also breaks the weak $\mathbb{Z}_{2}$ symmetry generated by the simultaneous bit-flip on both legs of the ladder. Consequently, the ground state of the combined Hamiltonian (16),(32) is not expected to exhibit SSB. We confirm this by calculating the Binder cumulant as a function of the transverse field using DMRG for various lengths of the spin ladder, see Fig. 7(a). Note that the DMRG calculation here is done for a non-Hermitian Hamiltonian and yet yields a real ground state energy and observables, thus manifesting the aforementioned $\mathcal{PT}$ symmetry. As opposed to the $U_{4}$ versus $g$ plots presented earlier (see Fig. 3), here the curves corresponding to different system sizes intersect at different $g$ values, demonstrating the absence of a scale invariant crossing, and, hence, the phase transition. This behavior of the Binder cumulant closely resembles the one occuring in the ferromagnetic mixed-field Ising chain, where it is well-known that the Ising symmetry and the SSB order are absent. Given these results, we conclude that the SSB ordered phase and phase transition present in the steady state of noiseless ITE under the ferromagnetic quantum Ising Hamiltonian does not survive when the ITE is subject to AD noise.
@@ IV Conclusion
In this paper we studied whether ITE-based preparation of SSB ground state order on a quantum computer is robust to noise. For the sake of generality and simplicity, we did not consider any particular ITE algorithm, instead focusing on a black box approach representing ITE as a Trotterized imaginary-time “circuit” and investigating the SSB in the mixed steady state of that circuit. The main conclusion of our work is that, given a target Hamiltonian with a ground state that spontaneously breaks a symmetry, the preparation of the ground state order using the ITE is robust against noise, as long as the noise channel respects the weak version of that symmetry. The strong version of the symmetry, on the other hand, does not play a role in the preparation of the SSB order, since all the relevant observables are linear in the density operator.
@@
To illustrate these conclusions, we employed the Choi-Jamiołkowski isomorphism and in the limit of weak noise and small imaginary-time Trotter step mapped the noisy ITE to a noiseless ITE under an effective Hamiltonian in the doubled Hilbert space. This effective Hamiltonian, in general, corresponds to two copies of the original target Hamiltonian coupled by noise-induced terms, and in the one-dimensional case its ground state can be efficiently computed using the DMRG. Focusing on a concrete example of the one-dimensional transverse-field Ising model as a target Hamiltonian and various types of noise affecting the ITE, we employed DMRG to compute the ground state of the effective Hamiltonian, which in this case corresponds to an Ising spin ladder, and applied the obtained results to study the SSB order and the associated phase transition in the mixed steady state of the noisy ITE. In addition, for comparatively small system sizes we directly simulated the evolution of the density matrix under the noisy ITE to show that our conclusions hold in the case of a small but finite Trotter step.
@@
We conclude by discussing implications of our work for practical simulations of the ITE on noisy quantum devices. While in this work we modeled the ITE as a black box, an algorithm implementing the ITE, which is a non-unitary protocol, on a device [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] requires the introduction of certain nontrivial operations, such as mid-circuit measurements and feedback [28]. Still, we generally expect our results to hold for a realistic protocol implementing the ITE as long as at each step the (noiseless) protocol well approximates an ITE step, and the state is moved closer to the target ground state. However, not all protocols implement the ITE in that fashion, and it would be interesting to “open up” the ITE black box and investigate the robustness of phase preparation against noise in various specific ITE quantum algorithms. Another potential future direction is to study how the Trotter error affects the preparation of the ground state order, and whether a phenomenon like the Trotter-error-induced phase transition 333for Trotter transitions in the context of quantum dynamics and information spreading, see Ref. [110, 111, 112, 113] can occur within this model.
Another interesting direction to explore within this context is the use of noise tailoring techniques like Pauli twirling [99, 100] and probabilistic error cancellation/reduction [101, 102, 103]. In this approach, the target circuit implementing, e.g., noisy ITE is replaced by a (quasi)probabilistic ensemble of circuits over which observable calculations are averaged. Up to finite sampling error and possible signal-to-noise issues, the result is an effective circuit with a modified noise. For example, Pauli twirling of Clifford gates can convert a generic noise channel into a Pauli noise channel [99]. This is potentially valuable in cases, like that of the ferromagnetic order in the Ising chain considered here, where it is desirable for the noise to take a Pauli form. The possible extension of our results to specific implementations of ITE quantum algorithms in concert with such noise tailoring and error mitigation techniques thus offers the intriguing prospect of a noise-resilient approach to quantum state preparation.

### v2

@@ I Introduction
Another conceptual approach to ground state preparation, which can potentially be more suitable for near-term implementation, is based on imaginary time evolution (ITE) [3, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Being a non-unitary operation, the ITE cannot be implemented using only unitary gates; nevertheless, various algorithms have been designed that effectively realize an approximate ITE process on a quantum device. These algorithms either (1) perform local tomography to determine a unitary such that the unitarily-evolved state approximates the ITE-evolved state [11, 12, 13, 14, 15, 16] (this is known as quantum imaginary time evolution, or QITE), (2) classically construct an adiabatic protocol that emulates the ITE [17] (this is known as adiabatic quantum imaginary time evolution, or A-QITE), (3) take advantage of the variational principle [18, 19, 20, 21, 22, 23, 24, 25] (this is known as variational quantum imaginary time evolution, or VQITE), (4) utilize ancilla(e) and measurements [26, 27, 28] (this is known as probabilistic imaginary time evolution, or PITE), or (5) ancilla(e), measurements, and feedback [29]. These and other sophisticated protocols become increasingly realizable on present-day and near-term devices due to rapid hardware advances, such as the ever more efficient implementation of mid-circuit measurements and the adoption of dynamical circuits [30, 31, 32].
@@
In this work we ask whether ITE-based ground state preparation is robust to various types of noise commonly present in experimental setups. To keep our considerations as general and as simple as possible, we refrain from directly modeling a particular algorithm implementing the ITE, and instead consider a black box realizing a Trotterized version of ITE  [^1] This Trotterized approach to ITE is also implemented directly in a variety of classical algorithms for quantum many-body physics, including time-evolving block decimation [64, 65] and various quantum Monte Carlo approaches [66, 67, 68, 69]. A “circuit” constructed in this way can be thought of as a toy model of an ITE protocol, which we further subject to noise (see Fig. 1). Within this model the ITE effectively realizes an approximate (due to the Trotter error) weak projection onto a target ground state, which may belong to a nontrivial phase, whereas noise tends to trivialize the state it acts upon. Here we analyze this competition between ITE and noise by investigating whether the mixed state resulting from the noisy ITE process belongs to the same nontrivial phase as the targeted pure ground state.
@@
As a concrete example, we focus on the one-dimensional transverse-field Ising model whose ground-state quantum phase transition is a prototypical example of long-range order and spontaneous symmetry breaking (SSB). We investigate whether this order, and the quantum critical point from which it emerges, persists in the mixed steady state of the ITE at finite noise strength. To this end, in the limit of infinitesimal imaginary Trotter step we take advantage of the Choi-Jamiołkowski isomorphism [35, 36] and map the noisy ITE of interest to a noiseless ITE under an effective Hamiltonian in a doubled Hilbert space. This allows us to extract the steady state properties of the noisy ITE from the ground state of the effective Hamiltonian, which can be efficiently calculated using the density matrix renormalization group (DMRG) [37, 38, 39]. To further corroborate our findings and investigate the effect of imaginary-time Trotter error, we also perform direct simulations of our noisy ITE “circuit” for comparatively small system sizes. We find that the ordered phase is stable to sufficiently weak noise as long as the noise does not explicitly break the weak version [40, 41] of the symmetry that is spontaneously broken at the ground-state quantum critical point.
@@ III.1.1 Single-qubit bit-flip noise
where in the second equality we have used the fact that $e^{\mu\sum_{i}\sigma^{x}_{i}\otimes\sigma^{x}_{i}}=\prod_{i}e^{\mu\sigma^{x}_{i}\otimes\sigma^{x}_{i}}=\prod_{i}[\cosh\mu+\sinh\mu\,\sigma^{x}_{i}\otimes\sigma^{x}_{i}]$ and have introduced a parameter $\mu=\tanh^{-1}[p/(1-p)]$, such that $e^{-\mu}\cosh\mu=1-p$ and $e^{-\mu}\sinh\mu=p$. This parameter is real and positive for $20<p<1/$ 2 and in the limit of weak noise, $p\ll 1$, becomes $\mu\sim p+O(p^{2})$. Given Eqs. (7) and (12), in the doubled space the steady state of the noisy ITE can be written, up to an overall normalization prefactor, as
@@
To showcase the above statements, we numerically investigate [50] the symmetry breaking order-disorder transition in the steady state of the noisy ITE by computing the ground state of the spin ladder Hamiltonian (16)-(17) using the DMRG [37, 38, 39] and then calculating the expectation values of observables in the steady state via Eq. (9). Numerical details of the DMRG calculations are presented in Appendix A. To identify the location of the transition as precisely as possible, we calculate the Binder cumulant [51, 52],
@@
Figure 3(a) depicts the Binder cumulant evaluated for the DMRG-calculated ground state of the spin ladder Hamiltonian (16)–(17) as a function of the transverse field $g$ for various values of the ladder length $L$ and open boundary conditions. The Ising couplings along the legs of the ladder are taken to be ferromagnetic, while the strength of the noise-induced couplings along the rungs is set to $\lambda=0.1$. For this value of $\lambda$ the critical point, identified through the scale invariance of the Binder cumulant, is located at $g_{c}\approx 0.94$, which is slightly lower than the noiseless value of $g_{c}=1$. The ordered phase is thus still present in the noisy model, although its extent in parameter space is smaller when compared to the noiseless case. The curves representing different system sizes collapse when the Binder cumulant is plotted as a function of $(g-g_{c})L$, which is illustrated by the inset in Fig. 3(a). The finite size scaling thus yields the correlation length critical exponent of $\nu=1$, which, as expected, belongs to the 2D Ising universality class. In Appendix B we also plot the dependence of the energy gap on $g-g_{c}$ near critical point, from which we extract dynamical critical exponent $z=1$, and showcase power-law decay of the $\langle\sigma^{z}_{i}\sigma^{z}_{j}\rangle$ correlation function at criticality, extracting critical exponent $\eta=1/4$. The values of these critical exponents are consistent with the 2D Ising universality class.
@@
We compute the two ground states, one with $\lambda=0.1$ and the other with $\lambda=0$, and the overlap between them using DMRG, and plot the resulting fidelity as a function of the transverse field in Fig. 4. For $g\ll 1$ ($g\gg 1$), both the noisy and the noiseless steady state are deep in their respective ordered (disordered) phases, and the fidelity between them is close to one. On the other hand, for values of $g$ in between the noiseless and the noisy critical points (marked by dotted lines in the figure), the noisy steady state is disordered, while the noiseless one is ordered, and the fidelity dips drastically toward zero. The dip is stronger for larger system sizes, with the fidelity reaching zero for $8002L=$ 800. We note that the fidelity also decreases with system size even deep in the ordered/disordered phase, albeit the decrease is much smaller there than in the region between the critical points. We attribute this decrease to the orthogonality catastrophe and expect the fidelity to decay to zero exponentially with system size for any value of $g$ in the thermodynamic limit, $L\to\infty$. While we observe exponential decay within the region between the noiseless and the noisy critical points, unfortunately, the available system sizes preclude us from definitively inferring exponential decay outside of this region. We demonstrate scaling of the fidelity with system size in Appendix C.
@@ III.2 Amplitude damping noise channel
Figure 7: Binder cumulant plotted as a function of the transverse field and calculated using the DMRG for (a) the ferromagnetic (FM, $J=+1$) and (b) the antiferromagnetic (AFM, $J=-1$) Ising spin ladder model (16) when the interleg coupling is induced by the amplitude damping (AD) noise given by Eq. (32) with $\lambda_{z}=\lambda_{+}=0.4$. The curves are plotted for various ladder lengths $L$ in order to identify the absence (in the FM case) or the presence (in the AFM case) of the scale invariant critical point $g_{c}$; the total number of sites in the DMRG calculation is equal to $2L$. The inset in (b) depicts the Binder cumulant as a function of $(g-g_{c})L$ and demonstrates the collapse of the curves corresponding to different system sizes. Open boundary conditions are assumed throughout the calculations in both (a) and (b).
@@
If the Ising ladder has ferromagnetic couplings along its legs, the Hamiltonian (32), in addition to breaking the strong $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry, also breaks the weak $\mathbb{Z}_{2}$ symmetry generated by the simultaneous bit-flip on both legs of the ladder. Consequently, the ground state of the combined Hamiltonian (16),(32) is not expected to exhibit SSB. We confirm this by calculating the Binder cumulant as a function of the transverse field using DMRG for various lengths of the spin ladder, see Fig. 7(a). Note that the DMRG calculation here is done for a non-Hermitian Hamiltonian and yet yields a real ground state energy and observables, thus manifesting the aforementioned $\mathcal{PT}$ symmetry. As opposed to the $U_{4}$ versus $g$ plots presented earlier (see Fig. 3), here the curves corresponding to different system sizes intersect at different $g$ values, demonstrating the absence of a scale invariant crossing, and, hence, the phase transition. This behavior of the Binder cumulant closely resembles the one occurring in the ferromagnetic mixed-field Ising chain, where it is well-known that the Ising symmetry and the SSB order are absent. Given these results, we conclude that the SSB ordered phase and phase transition present in the steady state of noiseless ITE under the ferromagnetic quantum Ising Hamiltonian does not survive when the ITE is subject to AD noise. In addition to computing $U_{4}$, we compute fidelity between the target ground state and the steady state of the noisy ITE in the presence of the AD noise using Eqs. (19)-(20). These results are presented in Appendix D.
@@ IV Conclusion
In this paper we studied whether ITE-based preparation of SSB ground state order on a quantum computer is robust to noise. For the sake of generality and simplicity, we did not consider any particular ITE algorithm, instead focusing on a black box approach representing ITE as a Trotterized imaginary-time “circuit” and investigating the SSB in the mixed steady state of that circuit. The main conclusion of our work is that, given a target Hamiltonian with a ground state that spontaneously breaks a symmetry, the preparation of the ground state order using the ITE is robust against noise, as long as the noise channel respects the weak version of that symmetry.
@@
To illustrate this conclusion, we employed the Choi-Jamiołkowski isomorphism and in the limit of weak noise and small imaginary-time Trotter step mapped the noisy ITE to a noiseless ITE under an effective Hamiltonian in the doubled Hilbert space. Focusing on a concrete example of the one-dimensional transverse-field Ising model as a target Hamiltonian and various types of noise affecting the ITE, we employed DMRG to compute the ground state of the effective Hamiltonian and applied the obtained results to study the SSB order and the associated phase transition in the mixed steady state of the noisy ITE. In addition, for comparatively small system sizes we directly simulated the evolution of the density matrix under the noisy ITE to show that our conclusions hold in the case of a small but finite Trotter step.
We emphasize that the resilience to noise of the ITE-based ground state phase preparation comes from the presence of the gap in the Liouvillian describing the ITE, or, equivalently, in the effective Hamiltonian of the system in the doubled Hilbert space. This gap protects the system from small perturbations that are caused by noise. Therefore, we conjecture that the ITE preparation of a certain phase of matter should be robust to noise as long as the gap in the Liouvillian is present for this phase of matter (as well as the required symmetries, if any, are not broken). This includes symmetry-protected topological phases (SPTs) and topologically ordered phases. However, we leave a detailed analysis of the effect of noise on the ITE preparation of SPTs and topologically ordered phases for a future study. In addition, we would like to emphasize that in this work we have only considered gapped phases of matter. Investigating preparation of ground state phases and its robustness to noise for gapless systems would be an interesting topic for future analysis. Here we only note that the gaplessness may alter the convergence time of the ITE, which indeed depends on the size of the gap. Another interesting topic for a future study is a comparison between the noisy ITE-based ground state phase preparation and noisy adiabatic/quantum phase estimation protocols.
@@
We conclude by discussing implications of our work for practical simulations of the ITE on noisy quantum devices. While in this work we modeled the ITE as a black box, an algorithm implementing the ITE, which is a non-unitary protocol, on a device [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 25, 26, 27, 28, 29] requires the introduction of certain nontrivial operations, such as mid-circuit measurements and feedback [29]. Still, we generally expect our results to hold for a realistic protocol implementing the ITE as long as at each step the (noiseless) protocol well approximates an ITE step, and the state is moved closer to the target ground state. However, not all protocols implement the ITE in that fashion, and it would be interesting to “open up” the ITE black box and investigate the robustness of phase preparation against noise in various specific ITE quantum algorithms. Another potential topic for future studies is to investigate how the Trotter error affects the preparation of the ground state order. In addition, an interesting direction to explore within this context is the use of noise tailoring techniques like Pauli twirling [59, 60] and probabilistic error cancellation/reduction [61, 62, 63]. The possible extension of our results to specific implementations of ITE quantum algorithms in concert with such noise tailoring and error mitigation techniques offers the intriguing prospect of a noise-resilient approach to quantum state preparation.